Determine the time in seconds taken for the temperature to fall from 68.0oC to 62.1oC. Measure the rate of fall at temperature 68.0oC and 62.1oC.
1
Expert's answer
2020-05-08T16:22:30-0400
"T-T_s=(T_o-T_s)e^{-kt}"

"62.1-0=(68-0)e^{-0.277t}"

"t=0.328\ s"